OpenStudy (anonymous):

In the following system of equations, what does x equal? 0x – 8y = -16 -8x + 2y = 36

OpenStudy (anonymous):

another zero!!!!!

OpenStudy (anonymous):

-8y = -16 y = -16/-8 y = 2 -8x+2(2) = 36 -8x+4 = 36 -8x = 36-4 x = 32/-8 x = -4

OpenStudy (anonymous):

@satellite u r wrong